# Construction of a Basis of Kähler differentials for generic plane branches

### v1

@@ 3.5.1. Definition and first properties of terminal families
We say that $\mathcal{T}_{l}=\{\Omega_{1},\ldots\Omega_{2{\nu}_{l}}\}\subset{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ is a terminal family of level $l$ if the following conditions hold:
@@
Condition (b) indicates that the sequence of multiplicities “does not go too far” and indeed that all of them are less than or equal to the multiplicity of $df_{l+1}$ which is a form of level $l$ since it belongs to $\hat{M}_{2{\nu}_{l}}$. Condition (c) implies, on one hand, that the leading variables are not greater than $\beta_{l+1}$ for $l<g$, and that if they are $\beta_{l+1}$, then they appear with degree 00; and on the other hand, that the $c$-sequences are required to be strictly increasing. The last property is inherent to the study the semi-module of contacts of holomorphic $1$-forms; notice that Expression (3) describes the semigroup of the curves in ${\mathcal{C}}$ but $f_{0}^{k_{0}}f_{1}^{k_{1}}\cdots f_{r}^{k_{r}}\in{\Delta}_{\beta_{r},\leq}^{k_{r}-1}\setminus{\Delta}_{\beta_{r},<}^{k_{r}-1}$ if $r\geq 1$ and $k_{r}>0$ and thus those products do not have strictly increasing $c$-sequences. Condition (c) is used in the inductive steps of our construction.
The family $\mathcal{T}_{0}=\{dx,df_{1}\}$ is terminal of level 00 because $\Omega_{1}\in\Delta_{n,<}$ and $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$. This is our starting point.
@@
In a terminal family ${\mathcal{T}}_{l}$ of level $l<g$, the multiplicity $\nu_{\mathcal{C}}(\Omega)$ of any $1$-form $\Omega\in\mathcal{T}_{l}$ belongs to $[\beta_{l+1}]_{e_{l}}$ or $[0]_{e_{l}}$ (because of their corresponding leading variables). This is a key property for our later arguments. More precisely, we have
  • If $\Omega$ is of level $l-1$, then $\nu_{\mathcal{C}}(\Omega)\in(e_{l})$;
@@
  • Let be $\Omega$ of level $l$. Then $\Omega\in\Delta_{\beta_{l+1},<}^{0}\Leftrightarrow\nu_{\mathcal{C}}(\Omega)\in[\beta_{l+1}]_{e_{l}}$
Hence, a terminal family of level $l<g$ “covers” all the congruences modulo $n$ in $[(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ (cf. Definition 3.6).
@@
We have $[\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}=[(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ for $l<g$.
@@
This holds because, on one hand, there are $2\nu_{l}$ congruences modulo $n$ in the set $(e_{l})\cup\{\beta_{l+1}+(e_{l})\}$ ($\nu_{l}$ from $(e_{l})$ and other $\nu_{l}$ from $\beta_{l+1}+(e_{l})$); and on the other hand, $[\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}$ has $2\nu_{l}$ elements by definition of terminal family. ∎
@@ 3.5.2. From level $l$ to level $l+1$
Suppose $\mathcal{T}_{l}=\{\Omega_{1},\ldots,\Omega_{2{\nu}_{l}}\}$ is a terminal family of level $l$ for $0\leq l<g$. Our aim is to construct a terminal family of level $l+1$ from $\mathcal{T}_{l}$. This requires several stages. The first one (stage 00) is the construction of the initial family $\mathcal{G}_{l+1}^{0}$ of stage 00 (Definition 3.5).
@@
for any $(a,j)\in{\mathfrak{S}}$, the set $[\nu_{\mathcal{C}}(\mathcal{G}_{l+1}^{0})]_{n}$ contains at most $\nu_{l+1}=n/e_{l+1}$ elements. Notice also that, since $m\overline{\beta}_{l+1}\not\in(e_{l})$ for any $1\leq m<n_{l+1}$, we have
@@
if $(a,j),(b,k)\in{\mathfrak{S}}$, $(a,j)\neq(b,k)$ and $\nu_{\mathcal{C}}(\Omega_{j})$ and $\nu_{\mathcal{C}}(\Omega_{k})$ belong both to $(e_{l})$ or both to $[\beta_{l+1}]_{e_{l}}$. In particular any element of $[(e_{l+1})]_{n}$ has at most two representatives in $(\nu_{\mathcal{C}}({\Omega}_{j}^{0}))_{j=1}^{2\nu_{l+1}}$. Hence, $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{j=1}^{2\nu_{l+1}}$ defines exactly the $\nu_{l+1}$ classes modulo $n$ in $[(e_{l+1})]_{n}$ and each class has two representatives in it.
@@
    $\begin{split}\{[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}:j\in{\mathcal{L}}_{s}\}=[(e_{l+1})]_{n}&\ \ \mathrm{and}\\ \{[\nu_{\mathcal{C}}(\Omega_{j}^{s})-{\mathfrak{n}}^{s}(\beta_{l+1})]_{n}:j\in{\mathcal{U}}_{s}\}=[(e_{l+1})]_{n}.&\end{split}$
    In particular $[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}$ if $j\neq k$ and both $j,k$ belong to ${\mathcal{L}}_{s}$ or both belong to ${\mathcal{U}}_{s}$.
@@ 3.5.3. From stage $s$ to $s+1$ inside level $l+1$
Assume that $[\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ for $j\neq k$, and that $\Omega^{s}_{k}\in\Delta_{{\mathfrak{n}}^{s}(\beta_{l+1}),<}$ and $\Omega^{s}_{j}\in\cup_{\beta^{\prime}<{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$. Let $(c_{{\mathfrak{n}}^{r}(\beta_{l+1})})_{r\geq 0}$ be the $c$-sequence of $\Omega^{s}_{k}$. Then:
@@
Suppose that $[\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}$ for $j<k$, and that $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ belong to $\tilde{\Delta}_{\beta_{l+1},\leq}$. Denote by $(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),j})_{r\geq 0}$ and $(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),k})_{r\geq 0}$ the $c$-sequences of $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ respectively. If $c_{{\mathfrak{n}}(\beta_{l+1}),j}\neq c_{{\mathfrak{n}}(\beta_{l+1}),k}$ and
@@
  • Let $1\leq j\leq 2\nu_{l}$ such that $\Omega_{j}\in\Delta_{\beta^{\prime},<}$ with $\beta^{\prime}=\beta_{l+1}$. Then $j+2(n_{l+1}-1)\nu_{l}\in{\mathcal{U}}_{1}$ and $j+2r\nu_{l}\in{\mathcal{L}}_{1}$ for $20\leq r\leq n_{l+1}-$ 2.
@@
Since any $[x]_{n}\in[(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{1\leq j\leq 2\nu_{l+1}}$ by Remark 3.13, we deduce ${\mathcal{U}}_{1}=\{1\leq r\leq 2\nu_{l+1}:\Omega_{r}^{1}\neq\Omega_{r}^{0}\}$.
@@
$[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{1})]_{n}$
if $j\neq j^{\prime}$ and $j,j^{\prime}\in{\mathcal{L}}_{s}$. Thus $\{[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}:j\in{\mathcal{L}}_{1}\}$ and $[(e_{l+1})]_{n}$ are sets of cardinal $\nu_{l+1}$ such that the former is included in the latter and as a consequence they coincide. Analogously, $[\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k^{\prime}}^{0})]_{n}$ holds if $k\neq k^{\prime}$ and $k,k^{\prime}\in{\mathcal{U}}_{1}$. This property together with equation (18) imply that $\{[\nu_{\mathcal{C}}(\Omega_{k}^{1})-{\mathfrak{n}}(\beta_{l+1})]_{n}:k\in{\mathcal{U}}_{1}\}$ is a set of $\nu_{l+1}$ elements contained in $[(e_{l+1})]_{n}$ and hence they coincide. ∎
@@
As $\mathcal{G}_{l+1}^{s}$ is good, $\Omega_{j}^{s}\in\cup_{\beta^{\prime}\leq{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$ for any $1\leq j\leq 2{\nu}_{l+1}$. Since ${\mathfrak{n}}^{s}(\beta_{l+1})\in(e_{l+1})$ by hypothesis, any element of $[(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{s}))_{1\leq j\leq 2\nu_{l+1}}$ by property (IV) for $s$. Thus ${\mathcal{L}}_{s+1}$ and ${\mathcal{U}}_{s+1}$ have both $\nu_{l+1}$ elements.
@@
By definition, the maximum ${\mathcal{C}}$-order of a $1$-form in a terminal family $\mathcal{T}_{l}$ of level $l$ is $\overline{\beta}_{l+1}$. Since a $1$-form $\Omega_{k}^{0}$ of stage 00 and level $l+1$ is of the form $f_{l+1}^{a}\Omega$ where $0\leq a<n_{l+1}$ and $\Omega\in\mathcal{T}_{l}$, we obtain
@@
Case $l+1=g$. Since $e_{g}=1$, we obtain good families ${\mathcal{G}}_{g}^{1},{\mathcal{G}}_{g}^{2},\ldots$ by iterative applications of Proposition 3.5. Let ${\mathcal{G}}_{g}^{s}$ be a good family. Condition (b) is trivially satisfied. Let $\mathcal{B}_{s}$ be the set of congruences $[x]_{n}$ modulo $n$ such that there exist $\Omega^{s}_{j},\Omega_{k}^{s}$ with $[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}=[x]_{n}$ and $j<k\leq n$. We claim that $\mathcal{G}_{g}^{s}$ is terminal if and only if $\mathcal{B}_{s}=\emptyset$. Let us show the non-trivial implication, assume $\mathcal{B}_{s}=\emptyset$. Since this implies
[…]
@@ 3.6. The terminal family of level $g$
Thus, beginning with the $1$-forms $\Omega_{1}=dx$ and $\Omega_{2}=df_{1}$ of level 00, we obtain a terminal family of $1$-forms $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{2n})$ of level $g$ by applying $g$ times Proposition 3.6.
@@
Remark 3.18.
Since $\nu_{\mathcal{C}}(f_{l})=\overline{\beta}_{l}$ for $1\leq l\leq g$, it is easy to build an algorithm that provides $\Lambda_{\Gamma}$, for generic $\Gamma\in\mathcal{C}$, and whose input is $n,\beta_{1},\ldots,\beta_{g}$, or equivalently $n,\overline{\beta}_{1},\ldots,\overline{\beta}_{g}$, by Remark 3.8.

### v3

@@ 3.5.1. Definition and first properties of terminal families
We say that $\mathcal{T}_{l}=(\Omega_{1},\ldots\Omega_{2{\nu}_{l}})\subset{\Omega}_{\mathcal{C}}(\mathbb{C}^{2},0)$ is a terminal family of level $l$ if the following conditions hold:
@@
Condition (b) indicates that the sequence of multiplicities “does not go too far” and indeed that all of them are less than or equal to the multiplicity of $df_{l+1}$ which is a form of level $l$ since it belongs to $\hat{M}_{2{\nu}_{l}}$. Condition (c) implies, on one hand, that the leading variables are not greater than $\beta_{l+1}$ for $l<g$, and that if they are $\beta_{l+1}$, then they appear with degree 0; and on the other hand, that the $c$-sequences are required to be strictly increasing. The last property is inherent to the study the semi-module of contacts of holomorphic $1$-forms; notice that Expression (3) describes the semigroup of the curves in $\mathcal{C}$} but $f_{0}^{k_{0}}f_{1}^{k_{1}}\cdots f_{r}^{k_{r}}\in{\Delta}_{\beta_{r},\leq}^{k_{r}-1}\setminus{\Delta}_{\beta_{r},<}^{k_{r}-1}$ if $r\geq 1$ and $k_{r}>0$ and thus those products do not have strictly increasing $c$-sequences. Condition (c) is used in the inductive steps of our construction.
The family $\mathcal{T}_{0}=(dx,df_{1})$ is terminal of level 0 because $\Omega_{1}\in\Delta_{n,<}$ and $\Omega_{2}\in\Delta_{\beta_{1},<}^{0}$. This is our starting point.
@@
In a terminal family ${\mathcal{T}}_{l}$ of level $l<g$, the multiplicity $\nu_{\mathcal{C}}(\Omega)$ of any $1$-form $\Omega\in\mathcal{T}_{l}$ belongs to $\beta_{l+1}]_{e_{l}}$ or $0]_{e_{l}}$ (because of their corresponding leading variables). This is a key property for our later arguments. More precisely, we have
  • If $\Omega$ is of level l1l-1, then $\nu_{\mathcal{C}}(\Omega)\in(e_{l})$;
@@
  • Let be $\Omega$ of level $l$. Then $\Omega\in\Delta_{\beta_{l+1},<}^{0}\Leftrightarrow\nu_{\mathcal{C}}(\Omega)\in[\beta_{l+1}]_{e_{l}}$ .
Hence, a terminal family of level $l<g$ “covers” all the congruences modulo $n$ in $(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ (cf. Definition 3.6).
@@
We have $\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}=[(e_{l})]_{n}\cup[\beta_{l+1}+(e_{l})]_{n}$ for $l<g$.
@@
This holds because, on one hand, there are $2\nu_{l}$ congruences modulo $n$ in the set $(e_{l})\cup\{\beta_{l+1}+(e_{l})\}$ ($\nu_{l}$ from $(e_{l})$ and other $\nu_{l}$ from $\beta_{l+1}+(e_{l})$); and on the other hand, $\nu_{\mathcal{C}}(\mathcal{T}_{l})]_{n}$ has $2\nu_{l}$ elements by definition of terminal family. ∎
@@ 3.5.2. From level $l$ to level $l+1$
Suppose $\mathcal{T}_{l}=(\Omega_{1},\ldots,\Omega_{2{\nu}_{l}})$ is a terminal family of level $l$ for $0\leq l<g$. Our aim is to construct a terminal family of level $l+1$ from $\mathcal{T}_{l}$. This requires several stages. The first one (stage 0) is the construction of the initial family $\mathcal{G}_{l+1}^{0}$ of stage 0 (Definition 3.5).
@@
for any $(a,j)\in{\mathfrak{S}}$, the set $\nu_{\mathcal{C}}(\mathcal{G}_{l+1}^{0})]_{n}$ contains at most $\nu_{l+1}=n/e_{l+1}$ elements. Notice also that, since $m\overline{\beta}_{l+1}\not\in(e_{l})$ for any $1\leq m<n_{l+1}$, we have
@@
if $\in{\mathfrak{S}}, (a,j)(b,k)(a,j)\neq(b,k)$ and $\nu_{\mathcal{C}}(\Omega_{j})$ and $\nu_{\mathcal{C}}(\Omega_{k})$ belong both to $(e_{l})$ or both to $\beta_{l+1}]_{e_{l}}$. In particular any element of $(e_{l+1})]_{n}$ has at most two representatives in $(\nu_{\mathcal{C}}({\Omega}_{j}^{0}))_{j=1}^{2\nu_{l+1}}$. Hence, $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{j=1}^{2\nu_{l+1}}$ defines exactly the $\nu_{l+1}$ classes modulo $n$ in $(e_{l+1})]_{n}$ and each class has two representatives in it.
@@
[…]
    In particular $\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}$ if $j\neq k$ and both $j,k$ belong to ${\mathcal{L}}_{s}$ or both belong to ${\mathcal{U}}_{s}$.
@@ 3.5.3. From stage $s$ to $s+1$ inside level $l+1$
Assume that $\nu_{\mathcal{C}}(\Omega^{s}_{j})]_{n}=[\nu_{\mathcal{C}}(\Omega^{s}_{k})]_{n}$ for $j\neq k$, and that $\Omega^{s}_{k}\in\Delta_{{\mathfrak{n}}^{s}(\beta_{l+1}),<}$ and $\Omega^{s}_{j}\in\cup_{\beta^{\prime}<{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$. Let $(c_{{\mathfrak{n}}^{r}(\beta_{l+1})})_{r\geq 0}$ be the $c$-sequence of $\Omega^{s}_{k}$. Then:
@@
Suppose that $\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}$ for $j<k$, and that $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ belong to $\tilde{\Delta}_{\beta_{l+1},\leq}$. Denote by $(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),j})_{r\geq 0}$ and $(c_{{\mathfrak{n}}^{r}(\beta_{l+1}),k})_{r\geq 0}$ the $c$-sequences of $\Omega_{j}^{0}$ and $\Omega_{k}^{0}$ respectively. If $c_{{\mathfrak{n}}(\beta_{l+1}),j}\neq c_{{\mathfrak{n}}(\beta_{l+1}),k}$ and
@@
  • Let $1\leq j\leq 2\nu_{l}$ such that $\Omega_{j}\in\Delta_{\beta^{\prime},<}$ with $\beta^{\prime}=\beta_{l+1}$. Then $j+2(n_{l+1}-1)\nu_{l}\in{\mathcal{U}}_{1}$ and $j+2r\nu_{l}\in{\mathcal{L}}_{1}$ for $0\leq r\leq n_{l+1}-2$.
@@
Since any $x]_{n}\in[(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{0}))_{1\leq j\leq 2\nu_{l+1}}$ by Remark 3.13, we deduce ${\mathcal{U}}_{1}=\{1\leq r\leq 2\nu_{l+1}:\Omega_{r}^{1}\neq\Omega_{r}^{0}\}$.
@@
$\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{0})]_{n}=[\nu_{\mathcal{C}}(\Omega_{j^{\prime}}^{1})]_{n}$
if $j\neq j^{\prime}$ and $j,j^{\prime}\in{\mathcal{L}}_{s}$. Thus $\{[\nu_{\mathcal{C}}(\Omega_{j}^{1})]_{n}:j\in{\mathcal{L}}_{1}\}$ and $(e_{l+1})]_{n}$ are sets of cardinal $\nu_{l+1}$ such that the former is included in the latter and as a consequence they coincide. Analogously, $\nu_{\mathcal{C}}(\Omega_{k}^{0})]_{n}\neq[\nu_{\mathcal{C}}(\Omega_{k^{\prime}}^{0})]_{n}$ holds if $k\neq k^{\prime}$ and $k,k^{\prime}\in{\mathcal{U}}_{1}$. This property together with equation (18) imply that $\{[\nu_{\mathcal{C}}(\Omega_{k}^{1})-{\mathfrak{n}}(\beta_{l+1})]_{n}:k\in{\mathcal{U}}_{1}\}$ is a set of $\nu_{l+1}$ elements contained in $(e_{l+1})]_{n}$ and hence they coincide. ∎
@@
As $\mathcal{G}_{l+1}^{s}$ is good, $\Omega_{j}^{s}\in\cup_{\beta^{\prime}\leq{\mathfrak{n}}^{s}(\beta_{l+1})}\Delta_{\beta^{\prime},<}$ for any $1\leq j\leq 2{\nu}_{l+1}$. Since ${\mathfrak{n}}^{s}(\beta_{l+1})\in(e_{l+1})$ by hypothesis, any element of $(e_{l+1})]_{n}$ has exactly two representatives in $(\nu_{\mathcal{C}}(\Omega_{j}^{s}))_{1\leq j\leq 2\nu_{l+1}}$ by property (IV) for $s$. Thus ${\mathcal{L}}_{s+1}$ and ${\mathcal{U}}_{s+1}$ have both $\nu_{l+1}$ elements.
@@
By definition, the maximum $\mathcal{C}$}-order of a $1$-form in a terminal family $\mathcal{T}_{l}$ of level $l$ is $\overline{\beta}_{l+1}$. Since a $1$-form $\Omega_{k}^{0}$ of stage 0 and level $l+1$ is of the form $f_{l+1}^{a}\Omega$ where $0\leq a<n_{l+1}$ and $\Omega\in\mathcal{T}_{l}$, we obtain
@@
Case $l+1=g$. Since $e_{g}=1$, we obtain good families ${\mathcal{G}}_{g}^{1},{\mathcal{G}}_{g}^{2},\ldots$ by iterative applications of Proposition 3.5. Let ${\mathcal{G}}_{g}^{s}$ be a good family. Condition (b) is trivially satisfied. Let $\mathcal{B}_{s}$ be the set of congruences $x]_{n}$ modulo $n$ such that there exist $\Omega^{s}_{j},\Omega_{k}^{s}$ with $\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}=[\nu_{\mathcal{C}}(\Omega_{k}^{s})]_{n}=[x]_{n}$ and $j<k\leq n$. We claim that $\mathcal{G}_{g}^{s}$ is terminal if and only if $\mathcal{B}_{s}=\emptyset$. Let us show the non-trivial implication, assume $\mathcal{B}_{s}=\emptyset$. Since this implies
$\left\{[\nu_{\mathcal{C}}(\Omega_{j}^{s})]_{n}\ |\ 1\leq j\leq n\right\}=n,$
@@ 3.6. The terminal family of level $g$
Thus, beginning with the $1$-forms $\Omega_{1}=dx$ and $\Omega_{2}=df_{1}$ of level 0, we obtain a terminal family of $1$-forms $\mathbf{\Omega}=(\Omega_{1},\ldots,\Omega_{2n})$ of level $g$ by applying $g$ times Proposition 3.6. The next result implies Theorem 1.1 by Remark 3.1.